# Closure Discovery for Coarse-Grained Partial Differential Equations using Multi-Agent Reinforcement Learning

### v1

@@ Appendix A Adapted PPO Algorithm
We note that this notation contains the parameters $\phi$, which parameterize the underlying neural network.
@@
$L_{\Pi}(\boldsymbol{O}^{n},\boldsymbol{A}^{n},\theta_{k},\theta,\beta):=\frac{1}{\tilde{N}_{x}\cdot\tilde{N}_{y}}\sum_{i,j=1}^{\tilde{N}_{x},\tilde{N}_{y}}L_{\pi_{ij}}({O}^{n}_{ij},{A}^{n}_{ij},\theta_{k},\theta)-\beta H(\pi_{ij,\theta}),$
@@
The advantage estimates ${Adv}^{\pi_{ij,\theta_{k}}}$ are computed with generalized advantage estimation (GAE) (Schulman et al., 2018) using the output of the global value network $\mathcal{V}^{FCN}_{\phi}$.
@@
  Input: Initial policy parameters $\theta_{1}$, initial value function parameters $\phi_{1}$, clip ratio $\epsilon$, discount factor $\gamma$, entropy regularization weight $\beta$
@@
     Collect set of trajectories $D_{k}$ with the global policy $\Pi^{FCN}_{\theta_{k}}$ in the environment
@@
     Update $\theta$ by performing a SGD step on $\theta_{k+1}=\arg\max_{\theta}\frac{1}{\|D_{k}\|}\sum_{\boldsymbol{O}^{n},\boldsymbol{G}^{n}\in D_{k}}L_{\Pi}(\boldsymbol{O}^{n},\boldsymbol{A}^{n},\theta_{k},\theta,\beta)$
@@
Appendix B Additional Results
B.1 Advection Equation
Figure 7: $\boldsymbol{\psi}^{0}$ is sampled from the MNIST test set. The velocity field is sampled from $\mathcal{D}^{Vortex}_{Test}$ (See Section C.2.1). Here, the IC of the concentration comes from a different distribution than the one used for training.
Figure 8: $\boldsymbol{\psi}^{0}$ is a sample from the F-MNIST test set. The velocity field is sampled from $\mathcal{D}^{Vortex}_{Train}$ (See Section C.2.2). Note that the velocity field comes from a different distribution than the one used for training.
B.2 Burgers’ Equation
Figure 9: $\boldsymbol{\psi}^{0}$ is a sample from $\mathcal{D}_{Train}^{Vortex}$. We plot the velocity magnitude of every $20$th step of simulation with $100$ coarse time steps. The example shows that CNN-MARL does also qualitatively keep the simulation closer to the FGS. The failure of the CGS to account for the subgrid-scale dynamics leads to diverging trajectories.
Figure 10: Velocity magnitude plotted for a 100-step roll out. The set-up is the same as in Figure 9 and the example again shows, that CNN-MARL leads to qualitatively different dynamics during a roll-out that are closer to those of the FGS.
Appendix C Technical Details on Hyperparamters and Training Runs
During training, we use entropy regularization in the PPO objective with a factor of $0.1$ and $0.05$ for the advection and Burgers’ equation respectively to encourage exploration. The discount factor is set to $0.95$ and the learning rate to $1\cdot 10^{-5}$. Training is done over 2000 epochs. In each epoch, 1000 transitions are collected. One policy network update is performed after having collected one new episode. We use a batch size of 10 for training. The total number of trainable parameters amounts to $188,163$ and the entire training procedure took about 8 hours for the advection equation on an Nvidia A100 GPU. For the Burgers’ equation, training took about $30$ hours on the same hardware. We save the policy every 50 epochs and log the corresponding MAE between CGS and FGS after 50 time steps. For evaluation on the advection equation, we chose the policy from epoch 1500 because it had the lowest logged MAE value. Figure 11 and Figure 12 show the reward curves and evolutions of episode lengths. As expected, the episode length increases as the agents become better at keeping the CGS and FGS close to each other.
@@ Appendix C Technical Details on Hyperparamters and Training Runs
Table 3: Numerical values of the simulation parameters used for the advection CGS and FGS. Note that $\widetilde{\Delta t}$ is chosen to guarantee that the CFL condition in the CGS is fulfilled.
@@
Resulting other parameters:
@@
Table 4: Numerical values of the simulation parameters used for the Burgers’ CGS and FGS. Again, $\widetilde{\Delta t}$ is chosen to guarantee that the CFL condition in the CGS is fulfilled.
@@
Resulting other parameters:
@@
C.1 Receptive Field of FCN
In our CNN-MARL problem setting, the receptive field of the FCN corresponds to the observation $O_{ij}$ the agent at point $(i,j)$ is observing. In order to gain insight into this, we analyze the receptive field of our chosen architecture.
@@ C.1 Receptive Field of FCN
C.2 Diverse Velocity Field Generation
C.2.1 Distribution for Training
@@ C.2.2 Distribution for Testing
C.3 Computational Complexity
To quantitatively compare the execution times of the different simulations, we measure the runtime of performing one update step of the environment and report them in Table 6. As expected, CNN-MARL increases the runtime of the CGS. However, it stays below the FGS times by at least a factor of $5$. The difference is especially pronounced in the example of the advection equation, where the FGS uses a high order scheme on a fine grid, which leads to an execution time difference between CNN-MARL and FGS of more than an order of magnitude.
@@ C.3 Computational Complexity
CGS ACGS CNN-MARL FGS
Advection $\tiny 0.31\pm 0.00$ $\tiny 0.96\pm 0.00$ $\tiny 2.66\pm 0.96$ $\tiny 89.52\pm 0.47$
Burgers’ $\tiny 0.25\pm 0.00$ $\tiny 1.82\pm 0.01$ $\tiny 10.16\pm 0.03$
Appendix D Proof: Theoretically Optimal Action

### v3

@@ Appendix B Adapted PPO Algorithm
We note that this notation contains the weights $\phi$, which parameterize the underlying neural network.
@@
$L_{\Pi}(\boldsymbol{O}^{n},\boldsymbol{A}^{n},\theta_{k},\theta,\beta):=\frac{1}{\tilde{N}_{x}\cdot\tilde{N}_{y}}\sum_{i,j=1}^{\tilde{N}_{x},\tilde{N}_{y}}L_{\pi_{ij}}({\boldsymbol{O}}^{n}_{ij},\boldsymbol{A}^{n}_{ij},\theta_{k},\theta)-\beta H(\pi_{ij,\theta}),$
@@
The advantage estimates ${Adv}^{\pi_{ij,\theta_{k}}}$ are computed with generalized advantage estimation (GAE) [49] using the output of the global value network $\mathcal{V}^{FCN}_{\phi}$.
@@
  Input: Initial policy weights $\theta_{1}$, initial value function weights $\phi_{1}$, clip ratio $\epsilon$, discount factor $\gamma$, entropy regularization weight $\beta$
@@
     Collect set of trajectories $D_{k}$ with the global policy $\Pi^{FCN}_{\theta_{k}}$
@@
[…]
@@
Appendix C Additional Results
C.1 Advection Equation
Figure 6: $\boldsymbol{\psi}^{0}$ is sampled from the MNIST test set. The velocity field is sampled from $\mathcal{D}^{Vortex}_{Test}$ (See Section D.2.1). Here, the IC of the concentration comes from a different distribution than the one used for training.
Figure 7: $\boldsymbol{\psi}^{0}$ is a sample from the F-MNIST test set. The velocity field is sampled from $\mathcal{D}^{Vortex}_{Train}$ (See Section D.2.2). Note that the velocity field comes from a different distribution than the one used for training.
C.2 Burgers’ Equation
Figure 8: $\boldsymbol{\psi}^{0}$ is a sample from $\mathcal{D}_{Train}^{Vortex}$. We plot the velocity magnitude of every $20$th step of simulation with $100$ coarse time steps. The example shows that Closure-RL does also qualitatively keep the simulation closer to the FGS. The failure of the CGS to account for the subgrid-scale dynamics leads to diverging trajectories.
Figure 9: Velocity magnitude plotted for a 100-step roll out. The set-up is the same as in Figure 8 and the example again shows, that Closure-RL leads to qualitatively different dynamics during a roll-out that are closer to those of the FGS.
Appendix D Technical Details on Hyperparamters and Training Runs
During training, we use entropy regularization in the PPO objective with a factor of $0.1$ and $0.05$ for the advection and Burgers’ equation respectively to encourage exploration. The discount factor is set to $0.95$ and the learning rate to $1\cdot 10^{-5}$. Training is done over 2000 epochs. In each epoch, 1000 transitions are collected. One policy network update is performed after having collected one new episode. We use a batch size of 10 for training. The total number of trainable weights amounts to $188,163$ and the entire training procedure took about 8 hours for the advection equation on an Nvidia A100 GPU. For the Burgers’ equation, training took about $30$ hours on the same hardware. We save the policy every 50 epochs and log the corresponding MAE between CGS and FGS after 50 time steps. For evaluation on the advection equation, we chose the policy from epoch 1500 because it had the lowest logged MAE value. Figure 10 and Figure 11 show the reward curves and evolutions of episode lengths. As expected, the episode length increases as the agents become better at keeping the CGS and FGS close to each other.
@@ Appendix D Technical Details on Hyperparamters and Training Runs
Table 2: Numerical values used for the advection CGS and FGS. Note that $\widetilde{\Delta t}$ is chosen to guarantee that the CFL condition in the CGS is fulfilled.
@@
Resulting other values:
@@
Table 3: Numerical values used for the Burgers’ CGS and FGS. Again, $\widetilde{\Delta t}$ is chosen to guarantee that the CFL condition in the CGS is fulfilled.
@@
Resulting other values:
@@
D.1 Receptive Field of FCN
In our Closure-RL problem setting, the receptive field of the FCN corresponds to the observation $\boldsymbol{O}_{ij}$ the agent at point $(i,j)$ is observing. In order to gain insight into this, we analyze the receptive field of our chosen architecture.
@@ D.1 Receptive Field of FCN
D.2 Diverse Velocity Field Generation
D.2.1 Distribution for Training
@@ D.2.2 Distribution for Testing
D.3 Computational Complexity
To quantitatively compare the execution times of the different simulations, we measure the runtime of performing one update step of the environment and report them in Table 5. As expected, Closure-RL increases the runtime of the CGS. However, it stays below the FGS times by at least a factor of $5$. The difference is especially pronounced in the example of the advection equation, where the FGS uses a high order scheme on a fine grid, which leads to an execution time difference between Closure-RL and FGS of more than an order of magnitude.
@@ D.3 Computational Complexity
CGS ACGS Closure-RL FGS
Advection $0.31\pm 0.00$ $0.96\pm 0.00$ $2.66\pm 0.96$ $89.52\pm 0.47$
Burgers’ $0.25\pm 0.00$ $1.82\pm 0.01$ $10.16\pm 0.03$
We additionally profiled our code for the advection experiment and found that 71.3% of the runtime is spent on obtaining trajectories from the FGS simulation. Only 1.3 % of the time is spent on the CGS (excluding the forward pass through the FCN), highlighting the significant computational overhead of the finer grid. The forward passes through the neural network account for 5.0% of the time, while updating the policy takes another 5.5%. This is in agreement with the theoretical considerations above that show that the FGS is computationally more expensive than the CGS. We note that there is room for optimization of the numerical solution, as currently the generation of the FGS simulation is not parallelized.
Appendix E Addition of a Global Reward
To investigate the effect of adding a global problem specific reward to the reward function, we repeated our experiment for the advection equation and added a penalty based on the deviation from the conservation of the total mass to each local reward term. The deviation from the total mass was scaled so that both parts of the reward term had the same order of magnitude. The obtained mean absolute error reduction compared to CGS was approximately 52 percent after 50 time steps during predictions (the local-only reward resulted in 53 percent error reduction) and thus did not offer any additional benefits compared to our original formulation. The situation may require further investigation for different global constraints and different PDEs. However for the present study the local rewards seem to be sufficient in order to obtain good closures.
Appendix F Discussion of Architecture Choice
We have chosen the IRCNN architecture [48] to have a small receptive field such that our agents learn and act locally in a manner that is reminiscent of the numerical discretizations of PDEs. The IRCNN consists of multiple convolutional layers and has been shown to be very successful in image restoration tasks [9]. Architectures such as LoFi [50], which also only have a small receptive field, could be a possible architecture choice for Closure-RL as well, although this would require further investigations.
Other architectures such as U-Nets [51] or Vision-Transformers [52, 53, 54] are desired to capture multiscale features or long-range dependencies which would significantly increase the receptive field of each agent. Based on our studies regarding the interpretation of actions (see Appendix G) as well as the derivation of the theoretical optimal action for one of the cases explored in detail (see Section Appendix H), these dependencies are generally not needed, as local information should be sufficient to complement numerical schemes. We note that this is not necessarily the case for other systems of interest such as PDEs with forcing terms, PDEs with time-varying parameters or Integro-Differential Equations.
Appendix G Interpretation of Actions
We are able to derive the optimal update rule for the advection CGS that negates the errors introduced by the numerical scheme (see Appendix H for the derivation):
$\tilde{\epsilon}^{n}=\frac{\widetilde{\Delta t}}{2}\tilde{\psi}_{tt}^{n}+|\tilde{u}^{n}|\frac{\widetilde{\Delta x}}{2}\tilde{\psi}_{xx}^{n}+|\tilde{v}^{n}|\frac{\widetilde{\Delta y}}{2}\tilde{\psi}_{yy}^{n}+\mathcal{O}(\widetilde{\Delta t}^{2},\widetilde{\Delta x}^{2},\widetilde{\Delta y}^{2})$ (22)
In order to further quantify the similarity between the numerical estimates of $-\widetilde{\Delta t}\tilde{\epsilon}^{n-1}_{i,j}$ and the taken action, we compute the Pearson product-moment correlation coefficient for 100 samples. The results are presented in Table 6 and show that the learned actions as well as the optimal action of the CGS are highly correlated for all different combinations of seen and unseen ICs and PDEPs.
MNIST F-MNIST
$\mathcal{D}^{Vortex}_{Train}$ $0.82\pm 0.05$ $0.70\pm 0.08$
$\mathcal{D}^{Vortex}_{Test}$ $0.82\pm 0.03$ $0.72\pm 0.08$
Additionally, we note that the truncation error contains a second-order temporal derivative. At first glance, it might seem surprising that the model would be able to predict this temporal derivative as the agents can only observe the current time step. However, the observation contains both the PDE solution, i.e. the concentration for the present examples, as well as the PDEPs, i.e. the velocity fields. Thus, both the concentration and velocity field are passed into the FCN and due to the velocity fields enough information is present to infer this temporal derivative.
Appendix H Proof: Theoretically Optimal Action